Sol Summary – January 29th

Sol: 8
Summary Title: Extending the search
Author’s name: Andrew Wheeler
Mission Status: Nominal
Sol Activity Summary: Day dawned clear and cold (-5C) and augured well for the crew’s plans. Rob finally had both clear skies and a fully operational telescope to observe the sun. The resultant images are being processed and the astronomer’s report will have the details. The illnesses of yesterday are where they belong – yesterday. A localised morning EVA around the hab further tested the robot rover’s mobility and its ability to gather spectra whilst being remotely operated. After lunch, an EVA to the furthest north reached by the rovers at Tharsis Montes had a team searching for evaporites, concretions and petrified wood. Serendipitously, as the team was looking for rock markers on the top of Hab ridge, signs of live were discovered. Endoliths were observed beneath the surface of quartz cobbles. Three samples were selected for spectroscopic analysis and comparison to the vegetation spectra from the increasing MDRS-derived spectral library. The evening meal provided an opportunity to clear the shelves of opened foodstuffs with a simple meat and seven vegetables spread.
Look Ahead Plan: As the forecast predicts more clear and slightly warmer days, we intend to push as far north as the rovers’ limit will allow in our attempts to reach the white moon gypsum deposit to collect samples. A second EVA will examine anhydrite yardangs in an environmental degradation study.
Anomalies in work: Nothing to report.
Weather: Clear morning at -5 Celsius rising to mid to high teens during the day. No clouds and no precipitation. Humidity continuously dropping from near 60% to 30%. The wind was faint to gentle during the day.
Crew Physical Status: Nothing to report.
EVA: EVA 13 with Steve and Scott with robot rover trials at the Hab.
EVA 14 with Andrew, Clare and Alex to Tharsis Montes in search of evaporites, concretions and petrified wood.
Reports to be filed: Sol Summary, Commander’s Report, Greenhab Report, Operations Report (including an HSO report), EVA report, EVA Request(s), Astronomer report.
Support Requested: Nothing to request.

Operations Report – January 29th

SOL: 8
Name of person filing report: Scott Dorrington
Non-nominal systems: None
Notes on non-nominal systems: None
Spirit rover used:
Hours: 243.9
Beginning charge: 100%
Ending charge: 38%
Currently charging: No
Opportunity rover used:
Hours: 158.3
Beginning charge: 100%
Ending charge: 46%
Currently charging: No
Curiosity rover used: Not used
Perseverance rover used: Not used
General notes on rovers: None
Summary of Hab operations:
WATER USE: 26.17 gallons
Water (static tank): 275.53 gallons (21” depth)
Static tank pipe heater (on or off): On
Static tank heater (On or off): on
Toilet tank emptied: no
Summary of internet: Nominal during comms window
Summary of suits and radios:
Suit #8 still standing down pending new nuts in charging port. Suit #6 battery was more depleted than others. Charge was 6.75 Volts upon return. It was a long EVA (3 hrs) and the EVA location was rather dusty – could be contributing factors.
Summary of GreenHab operations: put notes here
WATER USE: 6 gallons
Heater: On automatic
Supplemental light: 10pm-2am
Harvest: 2x tomatoes (21 g)
Summary of ScienceDome operations: Not used
Dual split: Nominal
Summary of RAM operations: Not used
Summary of any observatory issues: See astronomy report
Summary of health and safety issues: None. Crew back to full health.
Questions, concerns and requests to Mission Support: None.

Journalist Report – January 29th

Staring at the Sun
Yesterday (Sol 7) marked the half-way point of our rotation. Aside from one easy EVA in the early-morning, and a jigsaw puzzle that tested our wits more than it probably should have (barely managed to complete it and sign the back), we all pretty much had the day off, and Andrew submitted his commander’s report in lieu of a journalist’s report.
That’s not to say it was all sunshine and rainbows – we had a few illnesses as well. Happy to say that as of writing this on Sol 8, everyone’s pretty much back to normal.
Today involved a pair of great EVAs. The first pitted Steve’s rover against the ‘Mars 500’ racetrack – i.e. a custom course designed to test how quickly the rover could traverse a set course while sampling terrain. Steve tried several runs – one involved driving blindly and being guided by Scott (something that might be necessary in an actual Martian environment), the next involved driving the rover directly (which understandably streamlined and sped up proceedings), and the last involved Steve driving the rover while collecting samples en route. He also did a human test to compare results (again, the human clocked in more impressive speeds). During this EVA, while not helping Steve, Scott also experimented some more with his angular measurements, using his theodolite app and new protractor techniques.
The second EVA explored the area west of Tharsis Montes (about 4km due north of the MDRS), in search of gypsum (the ‘drinkable rock’ that can be heated to yield water) and petrified wood. We found plenty of the former, not so much of the latter, and got some incredible views of Skyline Rim (the dried-up streambeds we were navigating were pretty spectacular as well). In practice, the search for gypsum felt almost a little too gamey – if it glinted in the sun, we picked it up. We also discovered some endolithic colonies (microbes that live beneath rocks) in the process, partially by accident.
Following Rob’s previous stint setting up the observatory’s telescope and studying the sun, today he managed to take a cool photo of the sun’s surface.
The crew also passed another cool milestone today… our first self-grown tomatoes! Scott came back from the Greenhab bearing gifts, aka the most ‘tomatoey’ tomatoes we’ve ever had. Hats off to Crew 290 for laying the groundwork – we hope to do the same for Crew 292.
Highlights of the day/yesterday: Rob’s pic of the sun, jigsaw sleuthing, tomatoes, hot beverages with dairy milk, Greenhab now has a bee.
Lowlights of the day: Illness, burned fritters, the saltiest chips imaginable.
– Alexander Tobal, Crew 291

GreenHab Report – January 29th

GreenHab Officer: Scott Dorrington
Environmental control: heater on automatic, door opened for 15 minutes at midday
Average temperatures:
08:43: 74.5 F, 18%
13:13: 102.2 F, 11% door open
13:30: 84.9 F, 11% door closed
14:17: 101.1 F, 12%
15:07: 105.4 F, 12%
16:00: 103.3 F, 20%
17:09: 89.1 F, 24%
17:46: 77.0 F, 27%
Hours of supplemental light: 4 hours overnight
Daily water usage for crops: 6 gallons
Daily water usage for research and/or other purposes: 0 gallons
Water in Blue Tank (200 gallon capacity): 118 gallons
Time(s) of watering for crops: 8:43 mist, 13:13 water + mist, 15:07 water, 16:00 mist, 17:45 a bit of water from vegetable cooking
Changes to crops:
Today was another sunny day, although temperatures were manageable. I checked in with the plants before my morning EVA to give them a mist. After returning around 1pm, I returned to give the plants their first watering of the day, and also to release another spider captured in the hab. I gave special care to the newly planted carrots, making sure to keep the soil nice and moist. I left the door open for around 15 minutes while watering, which brought the temperatures down from 102 F to 85 F. This was a strategic cooling to avoid high temperatures later in the day. It worked – temperatures rose again to around 100 and remained there for the rest of the day. At 3pm, I returned for the second watering of the day, and some shuffling of the plants to keep them out direct sunlight in the afternoon. On my way back from the GreenHab, I stopped in to the Observatory to get a closer look at the sun that has been keeping our plants alive (managed to see a prominence and a few sun spots). In the afternoon, I noticed the two red tomatoes on tomato #9 were looking very ripe. Today was harvest day. We brought them back to the hab and divided them up – 1/3 of a tomato each. Steve also ran them through his spectrifier before consumption, so we got to see some nice tomato spectral curves. The crew relished the fruits of our labour – we all commented on the very tomato-y flavour – much better than store bought. Certified fresh!
Harvest: Tomatoes x2 (21 g)
Support/supplies needed: None

EVA Report – January 29th

EVA #: 13
Author: Scott Dorrington & Steven Hobbs
Start time: 1000
End time: 1200
This EVA was conducted in the vicinity of the Hab, and had a dual-purpose. Scott walked 200 m away from the Hab to the mound at the intersection of Cow Dung road. This gave a good vantage point to a number of peaks in the area for wayfinding experiments, including Phobos Peak and the Beehive (described in EVA 5), North Ridge, Kissing Camel Ridge, and the Henry Mountains. Bearings were taken to each peak using a theodolite app to visually display azimuths. Sightings were also taken of the sun at various times through the EVA to be used as a reference point for calibrating these readings. Compass bearings were also taken for validation. Angle measurements were also taken between pairs of peaks using a handheld protractor. A different method was used, where the protractor was placed on the ground and arms aligned one at a time. These readings are expected to be more accurate than previous methods. Vertical angles were also taken from the base to peak of Phobos peak – this will be used to estimate distance from Phobos peak via trigonometric calculations. During these activities, a car was observed driving past along Cow Dung road.
Rover spectral sample EVA trials were conducted to characterise the utility of a Sojourner-sized rover in negotiating Martian terrain and capturing VIS/NIR spectra. Six samples collected from previous EVAs were set up within 40 m of the HAB. The Johan Rover was timed for its traversal to each site, while guided by a second astronaut. This simulated a remote operator being guided by an on-site remote sensor. A second run was conducted where the operator was placed in an elevated position to self-guide the robot, simulating control from an overwatch mission. A third run timed traverse and positioning the rover for spectral sampling while the final run was conducted by timing an astronaut traversing the distance. Six spectra were captured and compared with results gathered in a controlled environment.
Destination: Outside the Hab. 518150, 4250950 for rover tasks. Walking within 200m of the Hab to 518332 4250785 for wayfinding points.
Purpose of EVA: Continue collecting spectral data of regolith and minerals around the Hab. Additionally, collect bearings, angles, and theodolite data to various peaks for wayfinding research.
Coordinates (use UTM WSG 84): 518150, 4250950 for rover and 518332 4250785 for wayfinding
EVA Participants: Steven Hobbs, Scott Dorrington
Road(s) and routes per MDRS Map: Exit via Hab airlock and walk to the hill SSE to SE of the Hab, then walk to turn off to Cow Dung Rd 0110 for wayfinding points. Within 200m of the Hab walking distance.
Mode of travel: Walking only.
Vehicles you will be using (If applicable): n/a

Sol Summary – January 31st

Sol: 10
Summary Title: Crunching the numbers
Author’s name: Andrew Wheeler
Mission Status: Nominal
Sol Activity Summary: The mornings are on repeat; clear and cold (-4C) morning and perfect for the solar observatory. A very short duration, early morning EVA to in front of the science dome’s panorama windows saw the final re-deploy of the solar logger. Despite the longer nights, the battery has been keeping the sensors working longer than anticipated. The performance data is eagerly awaited. The remainder of the morning was spent cleaning, cataloging, photographing and identifying samples and then preparing some for both non-destructive and destructive analysis. Lunch was a repeat of the scones and soup mostly because we’re getting good at this. After lunch, some of the gypsum collected the day before from the White Moon locale that had been subjected to spectral analysis, was shown to be completely different to other gypsum spectra. The repeat analyses and comparisons will be interesting. Dinner, a layered meat, vegetable and potato bake based on Hunza pie awaits.
Look Ahead Plan: Final EVAs to locate points of geological interest for GPS recording and a continuation of data analysis from spectral and physical samples.
Anomalies in work: Back pack batteries – see operations report.
Weather: Clear morning at -4 Celsius rising to mid teens during the day. Altostratus clouds with cumulus clouds on the southern horizon building in the afternoon with no precipitation. Humidity continuously dropping from near 70% at sunrise to 30% in the afternoon. The wind was faint to gentle during the day.
Crew Physical Status: Head bump in the observatory – see operations report.
EVA: EVA 18 with Andrew and Steve to re-deploy the solar logger.
Reports to be filed: Sol Summary, Journalist’s Report, Greenhab Report, Operations Report (including an HSO report), EVA report, EVA Request(s), Astronomer’s report.
Support Requested: Nothing to request.

Operations Report – January 31st

SOL: 10
Name of person filing report: Scott Dorrington
Non-nominal systems: None
Notes on non-nominal systems: None
Spirit rover used: Not used
Opportunity rover used: Not used
Curiosity rover used: Not used
Perseverance rover used: Not used
General notes on rovers:
Summary of Hab operations:
WATER USE: 21.15 gallons
Water (static tank): 223.25 gallons (depth 25”)
Static tank pipe heater (on or off): On
Static tank heater (on or off): On
Toilet tank emptied: no
Summary of internet: Nominal during comms window
Summary of suits and radios:
Batteries on suits #6 and #7 were replaced with new ones. Old batteries were labelled and placed with other faulty batteries. Charged batteries for 1 hour after replacement, then ran fans for 1 hour, testing voltage before and after. The voltage on suit 7 is also now stable after discharging.
Suit 6: Start: 13.60 V, End: 10.78 V (after 1 hr running)
Suit 7: Start: 13.08 V, End: 12.38 V (after 1 hr running)
Suit 8: Used nut kit in the RAM to judge the size of the nut. Found it is bigger than M3 and smaller than M4. Likely an imperial measurement.
Suit 11: While inspecting other suits for charging port nuts, I pulled suit 11 off the bottom rack. One of the screws from the back fell to the floor (likely not secured). Opening the back, I discovered the battery cables were not attached to the battery terminals. Not sure if this is deliberate. Left as found. Replaced screws in back, half screwed in – not all the way tightened.
Summary of GreenHab operations:
WATER USE: 9 gallons
Heater: On
Supplemental light: 10pm – 2am
Harvest: None
Summary of ScienceDome operations: Andrew used Science Dome for processing of gypsum discovered during yesterday’s EVA to White Moon. Gypsum was crushed in mortal and pestle and heated on hotplate to extract water. Clare used Science Dome to wash rock samples from previous EVAs. Steve used Science Dome to gather spectral curves of rock samples. Scott observed activities in Science Dome.
Dual split: Nominal
Summary of RAM operations: Entered RAM on several occasions to retrieve and return tools. Used tin snips to cut up a can for gypsum processing. Used socket sets and nut kit to investigate size of nut used for charging port on one-piece EVA suits.
Summary of any observatory issues: See astronomy report
Summary of health and safety issues:
Crew member received graze on head from hitting it on the door frame of the observatory. Bandages applied. Monitoring.
Questions, concerns and requests to Mission Support:
Advise on state of suit #11. Should the battery terminals be attached and screws inserted.

Journalist Report – January 31st

Breaking out the Bubbles
Sol 10 was a fairly relaxed day. There was reduced focus on EVAs… with the crew having collected most of the samples they needed during earlier EVAs, today was more about processing what we already had.
For Andrew and Clare, this involved preparing their gypsum samples for water extraction. The samples first needed to be washed, to remove any lingering dirt and other contaminants, and then ground in a mortar and pestle. The resulting powder was then heated. This experiment, although small, yielded the amount of water expected and was generally a success.
At every step of the way, Steve also took the opportunity to spectrify the gypsum – in its dirty state, its clean state, its powdered state, and its anhydrite state (post-water extraction). This went really well, although the spectral readings of the powdered gypsum weren’t fully consistent with Steve’s previous readings of similar samples, possibly implying slight differences in chemical composition between them. The more excitable of us have named this strange new form of gypsum “Stevenite”.
Andrew and Clare also reacted their carbonate samples (or specifically, the clay matrices within the samples) with ascetic acid (vinegar) to extract carbon dioxide, resulting in some very visible effervescence. Meanwhile Rob spent some more valuable time in the observatory.
As the crew’s engineering officer, Scott alternated between spacesuit maintenance (several of the suits were having battery issues) and Greenhab duties. Apparently we have two cucumbers in the ‘soon-to-be-eaten’ pipeline, as well as a surprise new fledgling cucumber from one of the more feeble plants. We didn’t necessarily have high expectations of the cucumber plants given that they had a rough time prior to our rotation with heating issues, so as far as we’re concerned this is all a bonus – clearly they’re enjoying the steady warmth. If we don’t eat them, they’ll make a good welcome present for Crew 292.
Speaking of warmth, in half a week we’ll be flying back into an Australian summer. Alas, our time at the MDRS is steadily drawing to a close. We only have one more full day ‘in sim’ before the end of our rotation – as of 12:00 local time on 2 February, we’re free men and woman (or at least, able to leave the hab without gearing up and depressurising in the airlock). The plan is to make tomorrow count with several large EVAs.
Highlights of the day: Water from rock, discovering Stevenite, discovering dehydrated chicken, Scotty to Engineering, Rob’s epic ‘hunza’ pie, cucumber surprise, discussing volcanoes/space disasters/conspiracies, Clare’s ‘weird rock’ fizzed (in vinegar).
Lowlights of the day: Self-dissembling plant-beds, Clare apparently hates puns and the news as well, and also hasn’t watched any Futurama or Star Wars/Trek (and calls the Force “psychic powers”), spherical trigonometry.
– Alexander Tobal, Crew 291

Copyright © The Mars Society. All rights reserved. | Main Site